Starting from the equation of motion of a
linear harmonic oscillator, show that its
phase space is an ellipse.
1
Expert's answer
2019-05-03T09:25:43-0400

The equation of motion of a linear harmonic oscillator


"\ddot{x}(t)+\omega^2x(t)=0"

Solution of this equation defines the position of the oscillator with time

"x(t)=A\sin(\omega t+\phi)"

The linear momentum of a linear harmonic oscillator

"p(t)=m\dot{x}(t)=m\omega A\cos(\omega t+\phi)"

Using two last equations we obtain the phase trajectory equation

"\frac{p^2}{(m\omega A)^2}+\frac{x^2}{A^2}=1"

Therefore the phase space of a linear harmonic oscillator is an ellipse with semi-axes

"a=m\omega A, \quad b=A"